Theorems of addition of probability and theorems of multiplication of probabilities are used for solving
1. There are 30 students in the group. The marks for the control test were the
following: 8 students received “excellent” marks, 12 students - “good” marks, 8
students – “satisfactory”. What is the probability that 3 students invited for answer have “good” marks?
1
Expert's answer
2020-11-04T16:04:04-0500

P(3 students have "good") = P(1st has "good" AND 2nd has "good" AND 3rd has "good") =

= P(1st has "good") * P(2nd has "good") * P(3rd has "good") = "\frac{12}{30}\cdot\frac{11}{29}\cdot\frac{10}{28} = 0.05"